OpenStudy (idealist10):

Find the point(s) on the curve of 4x^2+9y^2=36 at which the curvature is largest.

OpenStudy (idealist10):

x^2/9+y^2/4=1 x^2/a^2+y^2/b^2=1 x(t)=acos(t) y(t)=bsin(t) x(t)=3cos(t) y(t)=2sin(t) ---------------------- \[k=\frac{ \left| x'y''-y'x'' \right| }{ [(x')^2+(y')^2]^{3/2} }\]

OpenStudy (idealist10):

x'(t)=-3sin(t) y'(t)=2cos(t) x"(t)=-3cos(t) y"(t)=-2sin(t) k=6/(5sin^2 t+4)^(3/2) (3cos(t), 2sin(t))=(x, y) sin^2 t=y^2/4 k=6/(5y^2/4+4)^(3/2) y^2/4=1-x^2/9 so \[k=\frac{ 162 }{ (81-5x^2)^{3/2} }\]

OpenStudy (idealist10):

Now I need to find the derivative of k, so \[k'=\frac{ 2430x(81-5x^2)^{1/2} }{ (81-5x^2)^{3} }\]

OpenStudy (idealist10):

Simplify, I got \[k'=\frac{ 2430x }{ (81-5x^2)^{5/2} }=0\]

OpenStudy (idealist10):

So 2430x=0 and x=0. So x is the critical point? So what's the correct answer?

OpenStudy (idealist10):

@myininaya @amistre64

myininaya (myininaya):

I think you might also have to consider the endpoints of the ellipse

OpenStudy (idealist10):

So how do I find the right answer?

OpenStudy (idealist10):

@SithsAndGiggles @thomaster @amistre64

OpenStudy (anonymous):

Using the formula you gave, you have \[k(t)=\frac{|x'y''-x''y'|}{\left(\left(x'\right)^2+\left(y'\right)^2\right)^{3/2}}\] with \(x(t)=3\cos t\) and \(y(t)=2\sin t\). Your derivatives are \[\begin{matrix}x'=-3\sin t&&x''=-3\cos t\\ y'=2\cos t&&y''=-2\sin t\end{matrix}\] So, \[k(t)=\frac{|6\sin^2t+6\cos^2t|}{\left(9\sin^2t+4\cos^2t\right)^{3/2}}=\frac{6}{\left(9\sin^2t+4\cos^2t\right)^{3/2}}\] The derivative of the curvature function is \[k'(t)=-\frac{90\cos t\sin t}{(4\cos^2t+9\sin^2t)^{5/2}}\] so you have critical points for the values of \(t\) that give \[\cos t\sin t=0\] which are easy to find.

OpenStudy (idealist10):

So t=pi/2, pi, 3pi/2, am I right?

OpenStudy (idealist10):

@myininaya

OpenStudy (anonymous):

And \(t=0\), or \(t=2\pi\). This means you have to check four points.

OpenStudy (idealist10):

How did you get t=0, 2pi? And what to do after finding t?

OpenStudy (anonymous):

When \(t=0\) you get the same point as when \(t=2\pi\) (since \(0\) and \(2\pi\) correspond to the same angle on the unit circle). You get these from the fact that \(\sin t=0\) for \(t=n\pi\). Now you check the values of the curvature for these \(t\). For example, \(k(0)=\dfrac{6}{4^{3/2}}=\dfrac{3}{4}\) and \(k\left(\dfrac{\pi}{2}\right)=\dfrac{6}{9^{3/2}}=\dfrac{2}{9}\). Clearly the curvature at \(t=0\) is greater than at \(t=\dfrac{\pi}{2}\). This result makes sense, since the ellipse drawn below bows out more on the major axis (the horizontal axis) than it does on the minor axis: